Suppose a company has an ROE of 28.6 percent, total asset turnover ratio of 1.4 and equity multiplier of 01.75. If the average profit margin for other firms in the same industry is 16.40 percent, how much is this firm operating at below industry standards?(Express as a percentage to the nearest hundredth)

What annual rate of return is earned on a $8,950 investment that grows to $19,765 in eight years?

A.10.05%

B.10.41%

C.11.09%

D.11.73%

E.12.29%

.Assume you borrow $500 from a payday lender. The terms are that you must pay a fee of $97 in advance (today) and one year from now you need to repay $785. What implied interest rate are you paying?

A.43.09%

B.57.00%

C.76.47%

D.81.23%

E.94.79%.A stock investor deposited $3,450 six years ago. Today the account is valued at $2,180. What annual rate of return has this investor earned?

A.–6.65%

B.–7.37%

C.–8.45%

D.–9.74%

E.less than –9.74%
Answer #1

(1) ROE = 28.6 %, Total Asset Turnover Ratio = 1.4 and Equity Multiplier = 1.75

Let the company profit margin be P

Therefore, 28.6 = P x 1.4 x 1.75

P = 28.6 / (1.75 x 1.4) = 11.6735 % , Industry Average Profit MArgin = 16.4 %

Difference = 16.4 - 11.6735 = 4.72653 %

NOTE: Please raise separate queries for solutions to the remaining unrelated questions, as one query is restricted to the solution of only one complete question with up to four sub-parts.